OpenStudy (anonymous):

I=PRT; I=7280, P=26,000, R=0.O7 (SIMPLE INTEREST FORMULAR.)

hero (hero):

Okay....did you find T yet?

OpenStudy (anonymous):

NOOOOOOO

hero (hero):

T = I/(PR)

hero (hero):

T = 7280/(26,000*.07)

OpenStudy (anonymous):

PLEASE IS DAT THE FINAL ANSWER?

OpenStudy (anonymous):

7280=26000*0.07*t t=7280/(26000*0.07)

hero (hero):

No. The answer has to be calculated

OpenStudy (anonymous):

HOW CAN U HELP ME HW TO CALCULATE IT PLS

hero (hero):

Well, there is something called a calculator believe it or not. Thank goodness we live in an age where we can make such calculations using computational devices.

OpenStudy (anonymous):

T=4

hero (hero):

If you don't understand what I wrote, you shouldn't be taking math

OpenStudy (anonymous):

WTF

hero (hero):

I wasn't saying that to be rude. I was being honest.

OpenStudy (anonymous):

Completely agree

OpenStudy (anonymous):

Hero, im rubbing off on you :)

OpenStudy (anonymous):

I GOT THE ANSWER NOW 4

hero (hero):

Really? I doubt it. This is the first person I've come across that didn't understand division, computational evaluation or calculator.

OpenStudy (anonymous):

I SWEAR THE ANSWER IS 4

OpenStudy (anonymous):

now you see why i respond the way i respond sometimes

hero (hero):

Well, I didn't ask what grade he or she was in, so I'll take that into consideration.

OpenStudy (anonymous):

right, after you said " you shouldnt be taking math"

hero (hero):

Yeah, I see why they created delete buttons. Just in case you say something you might want to take back.

OpenStudy (anonymous):

HEY I HAVE ANOTHER QUESTION CAN U HELP ME WITH IT?

OpenStudy (anonymous):

hmm thats called regret

hero (hero):

No it's called un-insightful assumption

OpenStudy (anonymous):

yeah, you say tomato ,i say to-mato

hero (hero):

lol

OpenStudy (anonymous):

substitute the given values into the given formular and solve for the unknown variable if neccessary round to one decimal place (use the approximation 3.14 for ^) C=2^r; C=18.8 (CIRCUMFERENCE OF A CIRCLE)

OpenStudy (anonymous):

\[C=2\pi*r\]\[r=C/(2\pi)\] r=18.8/(2*3.14)